This is page 448 Printer: Opaque this

Bibliography

Aberth, O., Iteration methods for finding all zeros of a polynomial simultaneously, Math.

Comp. 27 (1973), 339–344. Abramowitz, M., & I.A. Stegun, Handbook of Mathematical Functions, Dover, New York,

NY, 1965. Achiezer, N.I., Theory of Approximation, Frederick Ungar, New York, NY, 1956. Ahlfors, L.V., Complex Analysis, McGraw-Hill, New York, NY, 1953. Aho, A.V., J.E. Hopcroft, & J.D. Ullman, The Design and Analysis of Computer Algo-

rithms, Addison-Wesley, Reading, MA, 1974. Anderson, J.M., Müntz-Szász type approximation and the angular growth of lacunary

integral functions, Trans. Amer. Math. Soc. 169 (1972), 237–248. Anderson, N., E.B. Saff, & R.S. Varga, On the Eneström-Kakeya theorem and its sharp-

ness, Linear Algebra and Appl. 28 (1979), 5–16. Anderson, N., E.B. Saff, & R.S. Varga, An extension of the Eneström-Kakeya theorem

and its sharpness, SIAM J. Math. Anal. 28 (1981), 10–22. Ankeny, N.C., & T.J. Rivlin, On a theorem of Bernstein, Pacific J. Math. 5 (1955),

849–852. Arestov, V.V., On integral inequalities for trigonometric polynomials and their deriva-

tives, Izv. Akad. Nauk SSSR Ser. Mat. 45 (1981), 3–22. Ash, R.B., Complex Variables, Academic Press, New York, NY, 1971.

Bibliography 449

Askey, R., & M. Ismail, Recurrence Relations, Continued Fractions and Orthogonal Polynomials, vol. 300, Mem. Amer. Math. Soc., Amer. Math. Soc., Providence, RI, 1984.

Aumann, G., Satz über das Verhalten von Polynomen auf Kontinuen, Sitz. Preuss. Akad., Wiss. Phys-Math. Kl. (1933), 924–931.

Aziz, A., & Q.M. Dawood, Inequalities for a polynomial and its derivative, J. Approx. Theory 54 (1988), 306–313.

Badalyan, G.V., Generalization of Legendre polynomials and some of their applications, Akad. Nauk Armyan. SSR Izv. Ser. Fiz.-Mat. Estest. Tekhn. Nauk 8 (1955), 1–28 (in Russian, Armanian summary).

Baishanski, B., Approximation by polynomials of given length, Illinois J. Math. 27 (1983), 449–458.

Baishanski, B., & R. Bojanic, An estimate for the coefficients of polynomials of given length, J. Approx. Theory 24 (1980), 181–188.

Bak, J., & D.J. Newman, Rational combinations of x^{λ_{k}} , λ_{k} ≥ 0 are always dense in C[0, 1], J. Approx. Theory 23 (1978), 155–157.

Baker, G.A., & P. Graves Morris, Padé Approximants, Part I, Basic Theory, Part II, Extensions and Applications, Addison-Wesley, Reading, MA, 1981.

Barbeau, E.J., Polynomials, Springer-Verlag, New York, NY, 1989.

Barnard, R.W., W. Dayawansa, K. Pearce, & D. Weinberg, Polynomials with nonnegative coefficients, Proc. Amer. Math. Soc. 113 (1991), 77–85.

Beauzamy, B., E. Bombieri, P. Enflo, & H.L. Montgomery, Products of polynomials in many variables, J. Number Theory 36 (1990), 219–245.

Beauzamy, B., & P. Enflo, Estimations de produits de polynômes, J. Number Theory 21 (1985), 390–412.

Beauzamy, B., P. Enflo, & P. Wang, Quantitative estimates for polynomials in one or several variables: From analysis and number theory to symbolic and massively parallel computation, Math. Mag. 67 (1994), 243–257.

Bernstein, S.N., Sur l’ordre de la meilleure approximation des functions continues par des polynômes de degré donné, Memoires de l’Académie Royale de Belgique 4 (1912), 1–103.

Bernstein, S.N., Sur la meilleure approximation de |x| par des polynômes de degrés donnés, Acta. Math. 37 (1913), 1–57.

Bernstein, S.N., Sur la représentation des polynômes positifs, Soobshch. Kharkow matem. ob-va, series 2 14 (1915), 227–228.

Bernstein, S.N., Sur une propriété des fonctions entières, C. R. Acad. Sci. Paris 176 (1923), 1603–1605.

450 Bibliography

Bernstein, S.N., Collected Works: Vol. I, Constr. Theory of Functions (1905–1930),

English Translation, Atomic Energy Commision, Springfield, VA, 1958. Beukers, F., A note on the irrationality of ζ(2) and ζ(3), Bull. London Math. Soc. 11

(1979), 268–272. Beukers, F., The value of polylogarithms, Colloq. Math. Soc. János Bolyai 34 (1981),

219–228. Bini, D., & V. Pan, Numerical and Algebraic Computations with Matrices and Polyno-

mials, Birkhäuser, Boston, MA, 1992. Block, H.D., & H.P. Thielman, Commutative polynomials, Quart. J. Math. Oxford Ser.

(2) 2 (1951), 241–243. Bloom, T., A spanning set for C(I^{n}), Trans. Amer. Math. Soc. 321 (1990), 741–759. Boas, R.P., Entire Functions, Academic Press, New York, NY, 1954. Bohman, H., On approximation of continuous and of analytic functions, Ark. Mat. 2

(1952), 43–56. Bojanov, B.D., An extension of the Markov inequality, J. Approx. Theory 35 (1982a),

181–190. Bojanov, B.D., Proof of a conjecture of Erdős about the longest polynomial, Proc. Amer.

Math. Soc. 84 (1982b), 99–103. Borodin A., & I. Munro, The Computational Complexity of Algebraic and Numeric

Problems, American Elsevier, New York, NY, 1975. Borosh, I., C.K. Chui, & P.W. Smith, Best uniform approximation from a collection of

subspaces, Math. Z. 156 (1977), 13–18. Borwein, J.M., & P.B. Borwein, Pi and the AGM – A Study in Analytic Number Theory

and Computational Complexity, Wiley, New York, NY, 1987. Borwein, P.B., Inequalities and inverse theorems in restricted rational approximation

theory, Canad. J. Math. 32 (1980), 354–361. Borwein, P.B., The size of {x : r′_{n}/r_{n} ≥ 1} and lower bounds for ∥e^{−x} − r_{n}∥, J. Approx.

Theory 36 (1982), 73–80. Borwein, P.B., Rational approximations to Stieltjes transforms, Math. Scand. 53 (1983),

114–124. Borwein, P.B., Markov’s inequality for polynomials with real zeros, Proc. Amer. Math.

Soc. 93 (1985), 43–48. Borwein, P.B., Zeros of Chebyshev polynomials in Markov systems, J. Approx. Theory

63 (1990), 56–64. Borwein, P.B., On the irrationality of ∑ 1/(q^{n} + r), J. Number Theory 37 (1991a),

253–259.

Borwein, P.B., Variations on Müntz’s theme, Bull. Canad. Math. Soc. 34 (1991b), 305– 310.

Borwein, P.B., On the irrationality of certain series, Math. Proc. Cambridge Philos. Soc. 112 (1992), 141–146.

Borwein, P.B., Exact inequalities for the norms of factors of polynomials, Can. J. Math. 46 (1994), 687–698.

Borwein, P.B., The arc length of the lemniscate |p(z)| = 1, Proc. Amer. Math. Soc. 123 (1995), 797–799.

Borwein, P.B., W. Chen, & K. Dilcher, Zeros of iterated integrals of polynomials, Can. J. Math. 11 (1995), 65–87.

Borwein, P.B., W. Dykshoorn, T. Erdélyi, & J. Zhang, Orthogonality and irrationality (to appear).

Borwein, P.B., & T. Erdélyi, Notes on lacunary Müntz polynomials, Israel J. Math. 76 (1991), 183–192.

Borwein, P.B., & T. Erdélyi, Remez-, Nikolskii- and Markov-type inequalities for generalized nonnegative polynomials with restricted zeros, Constr. Approx. 8 (1992), 343–362.

Borwein, P.B., & T. Erdélyi, Lacunary Müntz systems, Proc. Edinburgh Math. Soc. 36 (1993), 361–374.

Borwein, P.B., & T. Erdélyi, Sharp Markov-Bernstein type inequalities for classes of polynomials with restricted zeros, Constr. Approx. 10 (1994), 411–425.

Borwein, P.B., & T. Erdélyi, Dense Markov spaces and unbounded Bernstein inequalities, J. Approx. Theory 81 (1995a), 66-77.

Borwein, P.B., & T. Erdélyi, Müntz spaces and Remez inequalities, Bull. Amer. Math. Soc. 32 (1995b), 38–42.

Borwein, P.B., & T. Erdélyi, Upper bounds for the derivative of exponential sums, Proc. Amer. Math. Soc. 123 (1995c), 1481–1486.

Borwein, P.B., & T. Erdélyi, Generalizations of Müntz’s theorem via a Remez-type inequality for Müntz spaces (to appear 1).

Borwein, P.B., & T. Erdélyi, Markov and Bernstein type inequalities in L_{p} for classes of polynomials with constraints, J. London Math. Soc. (to appear 2).

Borwein, P.B., & T. Erdélyi, Newman’s inequality for Müntz polynomials on positive intervals, J. Approx. Theory (to appear 3).

Borwein, P.B., & T. Erdélyi, Sharp extensions of Bernstein’s inequality to rational spaces, Mathematika (to appear 4).

Borwein, P.B., & T. Erdélyi, The full Müntz theorem in C[0, 1] and L_{1}[0, 1], J. London Math. Soc. (to appear 5).

Borwein P.B., & T. Erdélyi, The L_{p} Version of Newman’s Inequality for lacunary polynomials, Proc. A.M.S. (to appear 6).

Borwein, P.B., T. Erdélyi, & J. Zhang, Chebyshev polynomials and Markov-Bernstein type inequalities for rational spaces, J. London Math. Soc. 50 (1994a), 501–519.

Borwein, P.B., T. Erdélyi, & J. Zhang, Müntz systems and orthogonal Müntz-Legendre polynomials, Trans. Amer. Math. Soc. 342 (1994b), 523–542.

Borwein P.B., E.A. Rakhmanov, & E.B. Saff, Rational approximation with varying weights I, Constr. Approx. (to appear).

Borwein, P.B., & E.B. Saff, On the denseness of weighted incomplete approximations, in A.A. Gonchar & E.B. Saff, Eds., Progress in Approximation Theory, Springer, New York, NY (1992), 419–429.

Borwein, P.B., & B. Shekhtman, The density of rational functions: A counterexample to a conjecture of D.J. Newman, Constr. Approx. 9 (1993), 105–110.

Boyd, D.W., Two sharp inequalities for the norm of a factor of a polynomial, Mathematika 39 (1992), 341–349.

Boyd, D.W., Bounds for the height of a factor of a polynomial in terms of Bombieri’s norms: I. the largest factor, J. Symbolic Comput. 16 (1993a), 115–130.

Boyd, D.W., Bounds for the height of a factor of a polynomial in terms of Bombieri’s norms: II. The smallest factor, J. Symbolic Comput. 16 (1993b), 131–145.

Boyd, D.W., Large factors of small polynomials, Contemp. Math. 166 (1994a).

Boyd, D.W., Sharp inequalities for the product of polynomials, Bull. London Math. Soc. 26 (1994b).

Boyer, C.B., A History of Mathematics, Wiley, New York, NY, 1968.

Braess, D., Nonlinear Approximation Theory, Springer-Verlag, Berlin, 1986.

Brent, R., The parallel evaluation of general arithmetic expressions, J. Assoc. Comput. Mach. 21 (1974), 201–206.

Browder, A., On Bernstein’s inequality and the norm of Hermitian operators, Amer. Math. Monthly 78 (1971), 871–873.

Brown, G., & E. Hewitt, A class of positive trigonometric sums, Math. Ann. 268 (1984), 91–122.

de Bruijn, N.G., Inequalities concerning polynomials in the complex domain, Nederl. Akad. Wetensch. Proc. 50 (1947), 1265–1272; Indag. Math. 9 (1947), 591–598.

de Bruijn, N.G., & T.A. Springer, On the zeros of composition-polynomials, Nederl. Akad. Wetensch. Proc. 50 (1947), 895–903; Indag. Math. 9 (1947), 406–414.

Bultheel, A., P. González-Vera, H. Hendriksen, & O. Njåstad, Orthogonal rational functions similar to Szegö polynomials, in: Orthogonal Polynomials and their Applications (C. Brezinski, L. Gori, and A. Ronveaux, Eds.), Scientific Publishing Co., Singapore, 1991, pp. 195–204.

Bibliography 453

Burton, D.M., The History of Mathematics, Allyn and Bacon, Newton, MA, 1985. Cauchy, A.L., Exercises de Mathématique, Oeuvres (2), vol. 9, p. 122, 1829. Cheney, E.W., Introduction to Approximation Theory, McGraw-Hill, New York, NY,

1966. Chihara, T.S., An Introduction to Orthogonal Polynomials, Gordon and Breach, New

York, NY, 1978. Chudnovsky, D.V., & G.V. Chudnovsky, Padé and rational approximation to systems

of functions and their arithmetic applications, in: Number Theory, New York, 1982 Seminar, Lecture Notes in Mathematics, vol. 1052, Springer-Verlag, Berlin, 1984, pp. 37–84.

Clark, A., Elements of Abstract Algebra, Wadsworth, Belmont, CA, 1971. Clarkson, J.A., & P. Erdős, Approximation by polynomials, Duke Math. J. 10 (1943),

5–11. Cody, W.J., W. Fraser, & J.F. Hart, Rational Chebyshev approximation using linear

equations, Numer. Math. 12 (1968), 242–251. van der Corput, J.G., & G. Schaake, Ungleichungen für Polynome und trigonometrische

Polynome, Compositio Math. 2 (1935), 321–361. Craven, T., G. Csordas, & W. Smith, The zeros of derivatives of entire functions and

the Pólya-Wiman Conjecture, Ann. of Math. 125 (1987), 405–431. Davis, P.J., Interpolation and Approximation, Dover, New York, NY, 1975. Dennis, J.E., & R.B. Schnabel, Numerical Methods for Unconstrained Optimization and

Nonlinear Equations, Prentice-Hall, Englewood Cliffs, NJ, 1983. DeVore, R.A., & G.G. Lorentz, Constructive Approximation, Springer-Verlag, Berlin,

1993. Dieudonné, M.J., La Théorie Analytique des Polynômes d’une Variable, Gauthier-

Villars, Paris, 1938. Dilcher, K., Real Wronskian zeros of polynomials with nonreal zeros, J. Math. Anal.

Appl. 154 (1991), 164–183. Ditzian, Z., Multivariate Landau-Kolmogorov-type inequality, Math. Proc. Cambridge

Philos. Soc. 105 (1989), 335–350. Ditzian, Z., & V. Totik, Moduli of Smoothness, Springer-Verlag, Berlin, 1988. Dolženko, E.P., Estimates of derivatives of rational functions, Izv. Akad. Nauk SSSR

Ser. Mat. 27 (1963), 9–28. Duffin, R.J., & A.C. Schaeffer, Some inequalities concerning functions of exponential

type, Bull. Amer. Math. Soc. 80 (1937), 1053–1070.

Duffin, R.J., & A.C. Scheaffer, Some properties of functions of exponential type, Bull. Amer. Math. Soc. 44 (1938), 236–240.

Duffin, R.J., & A.C. Scheaffer, A refinement of an inequality of the brothers Markoff, Trans. Amer. Math. Soc. 50 (1941), 517–528.

Durand, E., Solutions Numériques des Équations Algébriques, Tome I: Equations du Type F (x) = 0: Racines d’un Polynôme, Masson, Paris, 1960.

Eneström, G., Härledning af en allmän formel för antalet pensionärer..., Öfv. af. Kungl. Vetenskaps-Akademiens Förhandlingar 50 (1893), 405–415.

Erdélyi, A., et al., Higher Transcendental Functions, Vols. I–III, McGraw-Hill, New York, NY, 1953.

Erdélyi, T., Markov type estimates for derivatives of polynomials of special type, Acta Math. Hung. 51 (1988), 421–436.

Erdélyi, T., Markov-type estimates for certain classes of constrained polynomials, Constr. Approx. 5 (1989a), 347–356.

Erdélyi, T., The Remez inequality on the size of polynomials, in: Approximation Theory VI, Vol. 1 (C. K. Chui, L. L. Schumaker, & J. D. Ward, Eds.), Academic Press, Boston, MA, 1989b, pp. 243–246.

Erdélyi, T., Weighted Markov-type estimates for the derivatives of constrained polynomials on [0, ∞), J. Approx. Theory 58 (1989c), 213–231.

Erdélyi, T., A Markov-type inequality for the derivatives of constrained polynomials, J. Approx. Theory 63 (1990a), 321–334.

Erdélyi, T., A sharp Remez inequality on the size of constrained polynomials, J. Approx. Theory 63 (1990b), 335–337.

Erdélyi, T., Bernstein and Markov type inequalities for generalized non-negative polynomials, Canad. J. Math. 43 (1991a), 495–505.

Erdélyi, T., Bernstein-type inequalities for the derivative of constrained polynomials, Proc. Amer. Math. Soc. 112 (1991b), 829–838.

Erdélyi, T., Nikolskii-type inequalities for generalized polynomials and zeros of orthogonal polynomials, J. Approx. Theory 67 (1991c), 80–92.

Erdélyi, T., Estimates for the Lorentz degree of polynomials, J. Approx. Theory 67 (1991d), 187–198.

Erdélyi, T., Remez-type inequalities on the size of generalized polynomials, J. London Math. Soc. 45 (1992a), 255–264.

Erdélyi, T., Weighted Markov and Bernstein type inequalities for generalized nonnegative polynomials, J. Approx. Theory 68 (1992b), 283–305.

Erdélyi, T., Remez-type inequalities and their applications, J. Comput. Appl. Math. 47 (1993), 167-209.

Erdélyi, T., X. Li, & E.B. Saff, Remez and Nikolskii type inequalities for logarithmic potentials, SIAM J. Math. Anal. 25 (1994), 365–383.

Erdélyi, T., A. Magnus, & P. Nevai, Generalized Jacobi weights, Christoffel functions and Jacobi polynomials, SIAM J. Math. Anal. 25 (1994), 602–614.

Erdélyi T., A. Máté, & P. Nevai, Inequalities for generalized non-negative polynomials, Constr. Approx. 8 (1992), 241–255.

Erdélyi, T., & P. Nevai, Generalized Jacobi weights, Christoffel functions and zeros of orthogonal polynomials, J. Approx. Theory 69 (1992), 111–132.

Erdélyi, T., & J. Szabados, On polynomials with positive coefficients, J. Approx. Theory 54 (1988), 107–122.

Erdélyi, T., & J. Szabados, Bernstein-type inequalities for a class of polynomials, Acta Math. Hungar. 52 (1989a), 237–251.

Erdélyi, T., & J. Szabados, On trigonometric polynomials with positive coefficients, Studia Sci. Math. Hungar. 24 (1989b), 71–91.

Erdős, P., An extremum-problem concerning trigonometric polynomials, Acta. Sci. Math. Szeged 9 (1939), 113–115.

Erdős, P., On extremal properties of the derivatives of polynomials, Ann. of Math. 2 (1940), 310–313.

Erdős, P., Extremal problems on polynomials, in Approximation Theory II, Academic Press (1976), New York, NY, 347–355.

Erdős, P., F. Herzog, & G. Piranian, Metric properties of polynomials, J. Analyse Math. 6 (1958), 125–148.

Erdős, P., & P. Turán, On the distribution of roots of polynomials, Ann. of Math. 51 (1950), 105–119.

Erőd, J., On the lower bound of the maximum of certain polynomials, Mat. Fiz. Lap. 46 (1939), 58–83 (in Hungarian).

Favard, J., Sur les polynômes de Tchebycheff, C. R. Acad. Sci., Paris 200 (1935), 2052– 2055.

Feinerman, R.P., & D.J. Newman, Polynomial Approximation, Williams and Wilkins, Baltimore, MD, 1976.

Filaseta, M., Irreducibility criteria for polynomials with non-negative coefficients, Can. J. Math. 40 (1988), 339–351.

Fischer, B., Chebyshev polynomials for disjoint compact sets, Constr. Approx. 8 (1992), 309–329.

Frappier, C., & Q.I. Rahman, On an inequality of S. Bernstein, Canad. J. Math. 34 (1982), 932–944.

456 Bibliography

Frappier, C., Q.I. Rahman, & St. Ruscheweyh, New inequalities for polynomials, Trans.

Amer. Math. Soc. 288 (1985), 65–99. Freud G., Orthogonal Polynomials, Pergamon Press, Oxford, 1971. Fuchs, W.H.J., On the closure of {e^{−t}t^{a_{ν}} }, Proc. Cambridge Phil. Soc. 42 (1946), 91–

105. Gasper, G., & M. Rahman, Basic Hypergeometric Series, Cambridge University Press,

Cambridge, 1990. Gauss, C.F., Werke, Göttingen, 12 vols., 1863-1929. Gel’fond, A.O., Transcendental and Algebraic Numbers, Dover, New York, NY, 1960. Gencev, T.G., Inequalities for asymmetric entire functions of exponential type, Dokl.

Akad. Nauk SSSR 241 (1978), 1261–1264 (in Russian); Soviet Math. Dokl. 19 (1978), 981–985 (English translation).

Giroux, A., & Q.I. Rahman, Inequalities for polynomials with a prescribed zero, Trans.

Amer. Math. Soc. 193 (1974), 67–98. Glesser, P., Nouvelle majoration de la norme des facteurs d’un polynôme, C.R. Math.

Rep. Acad. Sci. Canada 12 (1990), 224–228. von Golitschek, M., A short proof of Müntz’s theorem, J. Approx. Theory 39 (1983),

394–395. von Golitschek, M., & G.G. Lorentz, Bernstein inequalities in L_{p}, 0 < p ≤ ∞, Rocky

Mountain J. Math. 19 (1989), 145–156. von Golitschek, M., G.G. Lorentz, & Y. Makovoz, Asymptotics of weighted polynomials,

in A.A. Gonchar, & E.B. Saff, Eds., Progress in Approximation Theory, Springer, New York, NY (1992), 431–451.

Govil, N.K., On the derivative of a polynomial, Proc. Amer. Math. Soc. 41 (1973),

543–546. Grace, J.H., The zeros of a polynomial, Proc. Cambridge Philos. Soc. 11 (1902), 352–357. Granville, A., Bounding the coefficients of a divisor of a given polynomial, Monatsh.

Math. 109 (1990), 271–277. Hardy, G.H., & J.E. Littlewood, A further note on the converse of Abel’s theorem, Proc.

London Math. Soc. 25 (1926), 219–236. He, X., Weighted Polynomial Approximation and Zeros of Faber Polynomials, Ph.D.

Dissertion, University of South Florida, Tampa, 1991. He, X., & X. Li, Uniform convergence of polynomials associated with varying Jocobi

weights, J. Math 21 (1991), 281–300. Henrici, P., Applied and Computational Complex Analysis, Vol. I, Wiley, New York,

NY, 1974.

Hille, E., G., Szegő, & J.D. Tamarkin, On some generalizations of a theorem of A. Markov, Duke Math. J. 3 (1937), 729–739.

Hölder, O, Über einen Mittelwertsatz, Göttinger Nachrichten (1889), 38–47.

Householder, A.S., The Numerical Treatment of a Single Nonlinear Equation, McGraw-Hill, New York, NY, 1970.

Ivanov, K.G., & V. Totik, Fast decreasing polynomials, Constr. Approx. 6 (1990), 1–20.

Ivanov, V.I., Certain inequalities for trigonometric polynomials and their derivatives in various metrics, Mat. Zametki 18 (1975), 489–498; Translation in Math. Notes 18 (1975), 880–885.

Jensen, J.L.W., Recherches sur la théorie des équations, Acta. Math. 36 (1913), 181–185.

Kakeya, S., On the limits of the roots of an algebraic equation with positive coefficients, Tôhoku Math. J. 2 (1912), 140–142.

Karlin, S., Total Positivity, Stanford University Press, Stanford, CA, 1968.

Karlin, S., & W.J. Studden, Tchebycheff Systems with Applications in Analysis and Statistics, Wiley, New York, NY, 1966.

Karp, R., & V. Ramachandran, Parallel algorithms for shared memory machines, in: Handbook of Theoretical Computer Science, Vol. A, Elsevier, Amsterdam, 1990, pp. 869–941.

Kemperman, J.H.B., & G.G. Lorentz, Bounds for polynomials with applications, Proc. Kon. Nederl. Akad. Wetensch. A 82 (1979), 13–26.

Kerner, I.O., Ein Gesamtschrittverfahren zur Berechnung der nullstellen von Polynomen, Numer. Math. 8 (1966), 290–294.

Khalilova, B.A., On some estimates for polynomials, Izv. Akad. Nauk Azerbaĭdzhan. SSSR Ser. Fiz.-Tekhn. Mat. Nauk 2 (1974), 46–55 (in Russian).

Klein, G., On the approximation of functions by polynomials (1951), Ph.D. Dissertation, University of Chicago, IL.

Kneser, H., Das Maximum des Produkts zweier Polynome, Sitz. Preuss. Akad. Wiss., Phys. - Math. Kl. (1934), 429–431.

Knuth, D.E., The Art of Computer Programming. Vol. 2: Seminumerical Algorithms, Addison-Wesley, Reading, MA, 1981.

Kolmogorov, A.N., On inequalities between the upper bounds of successive derivatives of an arbitrary function on an infinite interval, Amer. Math. Soc. Transl. Series 1, 2 (1962), 233–243.

Konjagin, S.V., Bounds on the derivatives of polynomials, Dokl. Akad. Nauk SSSR 243 (1978), 1116–1118 (in Russian); English translation: Soviet Math. Dokl. 19 (1978), 1477-1480.

458 Bibliography

Korovkin, P., On convergence of linear positive operators in the space of continuous

functions, Doklady Akad. Nauk SSSR 90 (1953), 961–964. Kristiansen, G.K., Proof of a polynomial conjecture, Proc. Amer. Math. Soc. 44 (1974),

58–60. Kristiansen, G.K., Some inequalities for algebraic and trigonometric polynomials, J.

London Math. Soc. 20 (1979), 300–314. Kristiansen, G.K., Solution to Problem 80 − 16, Siam Review 24 (1982), 77–78. Kroó, A., & F. Peherstorfer, On the distribution of extremal points of general Chebyshev

polynomials, Trans. Amer. Math. Soc. 329 (1992), 117–130. Kroó, A., & J.J. Swetits, On density of interpolation points a Kadec-type theorem and

Saff’s principle of contamination in L_{p} approximation, Constr. Approx. 8 (1992), 87–103.

Kroó, A., & J. Szabados, Constructive properties of self-reciprocal polynomials, Analysis

14 (1994a), 319–339. Kroó, A., & J. Szabados, Müntz-type problems for Bernstein polynomials, J. Approx.

Theory 78 (1994b), 446–457. Lachance, M.A., Bernstein and Markov inequalities for constrained polynomials, in: Ra-

tional Approximation and Interpolation, Lecture Notes in Mathematics, vol. 1045, Springer-Verlag, 1984, pp. 125–135.

Lachance, M.A., E.B. Saff, & R.S. Varga, Inequalities for polynomials with a prescribed

zero, Math. Z. 168 (1979), 105–116. Lax, P.D., Proof of a conjecture of P. Erdős on the derivative of a polynomial, Bull.

Amer. Math. Soc. 50 (1944), 509–513. Lebesgue, H., Sur l’approximation des fonctions, Bull. Sci. Math. 22 (1898), 278–287. Leviatan, D., Improved estimates in Müntz-Jackson theorems, in: Progress in Approxi-

mation Theory (P. Nevai & A. Pinkus, Eds.), Academic Press, 1991, pp. 575–582. Levin, A.L., & D.S. Lubinsky, Canonical Products and the weights exp(−|x|^{α}), α > 1,

with applications, J. Approx. Theory 49 (1987a), 149–169. Levin, A.L., & D.S. Lubinsky, Weights on the real line that admit good relative polyno-

mial approximation, with applications, J. Approx. Theory 49 (1987b), 170–195. Loomis, L.H., A note on the Hilbert transform, Bull. Amer. Math. Soc. 50 (1946), 1082–

1086. Lorentz, G.G., The degree of approximation by polynomials with positive coefficients,

Math. Ann. 151 (1963), 239–251. Lorentz, G.G., Approximation of Functions, 2nd ed., Chelsea, New York, NY, 1986a. Lorentz, G.G., Bernstein Polynomials, 2nd ed., Chelsea, New York, NY, 1986b.

Lorentz, G.G., Notes on Approximation, J. Approx. Theory 56 (1989), 360–365.

Lorentz, G.G., K. Jetter, & S.D. Riemenschneider, Birkhoff Interpolation, Addison-Wesley, Reading, MA, 1983.

Lorentz, G.G., M. von Golitschek, & Y. Makovoz, Constructive Approximation: Advanced Problems (to appear).

Lubinsky, D.S., Strong Asymptotics for Extremal Errors and Polynomials Associated with Erdős Weights, Pitman Research Notes in Mathematics Series, vol. 202, Longman House, Burnt Mill, Harlow, Essex, UK, 1989.

Lubinsky, D.S., & P. Nevai, Markov-Bernstein inequalities revisited, Approx. Theory Appl. 3 (1987), 98–119.

Lubinsky, D.S., & E.B. Saff, Strong asymptotics for extremal polynomials associated with weights on ℝ, Lecture Notes in Mathematics, vol. 1305, Springer-Verlag, Berlin, 1988a.

Lubinsky, D.S., & E.B. Saff, Uniform and mean approximation by certain weighted polynomials, with applications, Constr. Approx. 4 (1988b), 21–64.

Lubinsky, D.S., & E.B. Saff, Szegő asymptotics for non-Szegő weights on [−1, 1], in: Approximation Theory VI, Vol. 2 (C.K. Chui, L.L. Schumaker, & J.D. Ward, Eds.), Academic Press, Boston, MA, 1989, pp. 409–412.

Lubinsky, D.S., & E.B. Saff, Markov-Bernstein and Nikolskii inequalities, and Christoffel functions for exponential weights on [−1, 1], SIAM J. Math. Anal. 24 (1993), 528– 556.

Lubinsky, D.S., & Z. Ziegler, Coefficient bounds in the Lorentz representation of a polynomial, Canad. Math. Bull. 32 (1990), 197–206.

Lucas, F., Propriétés géométriques des fractions rationelles, C. R. Acad. Sci., Paris 77 (1874), 631–633.

Luxemburg, W.A.J., & J. Korevaar, Entire functions and Müntz-Szász type approximation, Trans. Amer. Math. Soc. 157 (1971), 23–37.

Mahler, K., Zur Approximation der Exponentialfunktion und des Logarithmus, J. Reine Angew. Math. 166 (1931), 118–150.

Mahler, K., An application of Jensen’s formula to polynomials, Mathematika 7 (1960), 98–100.

Mahler, K., On some inequalities for polynomials in several variables, J. London Math. Soc. 37 (1962), 341–344.

Mahler, K., A remark on a paper of mine on polynomials, Illinois J. Math. 8 (1964), 1–4.

Mairhuber, J., On Haar’s theorem concerning Chebyshev approximation problems having unique solutions, Proc. Amer. Math. Soc. 7 (1956), 609–615.

Malik, M.A., On the derivative of a polynomial, J. London Math. Soc. 8 (1969), 57–60.

Marden, M., Geometry of Polynomials, Math. Survey Monographs, No. 3, Amer. Math. Soc., Providence, RI, 1966.

Markov, A.A., Sur une question posée par D.I. Mendeleieff, Izv. Acad. Nauk St. Petersburg 62 (1889), 1–24.

Markov, V.A., Über Polynome die in einen gegebenen Intervalle möglichst wenig von Null abweichen, Math. Ann. 77 (1916), 213–258; the original appeared in Russian in 1892.

Máté, A., Inequalities for derivatives of polynomials with restricted zeros, Proc. Amer. Math. Soc. 88 (1981), 221–224.

Máté, A., & P. Nevai, Bernstein inequality in L_{p} for 0 < p < 1, and (C, 1) bounds of orthogonal polynomials, Ann. of Math. 111 (1980), 145–154.

Máté, A., P. Nevai, & V. Totik, Asymptotics for the ratio of leading coefficients of orthogonal polynomials on the unit circle, Constr. Approx. 1 (1985), 63–69.

Máté, A., P. Nevai, & W. van Assche, The support of measures associated with orthogonal polynomials and the spectra of the real self–adjoint operators, Rocky Mountain J. Math. 21 (1991), 501–527.

McCarthy, P.C., J.E. Sayre, & B.L.R. Shawyer, Generalized Legendre polynomials, J. Math. Anal. Appl. 177 (1993), 530–537.

Mhaskar, M.N., & E.B. Saff, Where does the sup norm of a weighted polynomial live?, Constr. Approx. 1 (1985), 71–91.

Mignotte, M., Some useful bounds, in: Computer Algebra: Symbolic and Algebraic Computation (Burchberger, B., et al., Eds.) Springer, New York, NY, 1982, pp. 259–263.

Mignotte, M., Mathematics for Computer Algebra, Spinger-Verlag, New York, NY, 1992.

Miller, M., Maximal polynomials and the Ilieff-Sendov conjecture, Trans. Amer. Math. Soc. 321 (1990), 285–303.

Milovanović, G.V., D.S. Mitrinović, & Th.M. Rassias, Topics in Polynomials: Extremal Problems, Inequalities, Zeros, World Scientific, Singapore, 1994.

Mitrinović, D.S., J.E. Pecaric, & A.M. Fink, Classical and New Inequalities in Analysis, Kluwer, Dordrecht, 1993.

Müntz, C., Über den Approximationssatz von Weierstrass, H.A. Schwarz Festschrift, Berlin (1914).

Natanson, I.P., Constructive Function Theory, Vol. 1, Frederick Ungar, New York, NY, 1964.

Nevai, P., Mean convergence of Lagrange interpolation, I, J. Approx. Theory 18 (1976), 363–377.

Bibliography 461

Nevai, P., Bernstein’s inequality in L_{p}, 0 < p < 1, J. Approx. Theory 27 (1979a), 239–

243. Nevai, P., Orthogonal Polynomials, vol. 213, Mem. Amer. Math. Soc., Amer. Math. Soc.,

Providence, RI, 1979b. Nevai, P., Géza Freud, Orthogonal polynomials and Christoffel functions. A case study,

J. Approx. Theory 48 (1986), 3–167. Nevai, P., Weakly convergent sequences of functions and orthogonal polynomials, J. Ap-

prox. Theory 65 (1991), 332–340. Nevai, P., & V. Totik, Weighted polynomial inequalities, Constr. Approx. 2 (1986), 113–

127. Nevai, P., & V. Totik, Sharp Nikolskii inequalities with exponential weights, Anal. Math.

13 (1987), 261–267. Newman, D.J., Derivative bounds for Müntz polynomials, J. Approx. Theory 18 (1976),

360–362. Newman, D.J., Approximation with Rational Functions, CBMS Regional Conf. Ser. in

Math. 41, Conf. Board Math. Sci., Washington, DC, 1978. Newman, D.J., & T.J. Rivlin, On polynomials with curved majorants, Canad. J. Math.

34 (1982), 961–968. Nürnberger, G., Approximation by Spline Functions, Springer-Verlag, Berlin, 1989. Ogawa, S., & K. Kitahara, An extension of Müntz’s theorem in multivariables, Bull.

Austral. Math. Soc. 36 (1987), 375–387. Olivier, P.F., & A.O. Watt, Polynomials with a prescribed zero and the Bernstein’s

inequality, Canad. J. Math. 45 (1993), 627–637. Operstein, V., Full Müntz theorem in L_{p}[0, 1], J. Approx. Theory (to appear). Osval’d, P., Some inequalities for trigonometric polynomials in L_{p}, 0 < p < 1, Izv.

Vyssh. Uchebn. Zaved. Mat. (7) 20 (1976), 65–75. Pan, V., Complexity of computations with matrices and polynomials, SIAM Rev. 34

(1992), 225–262. Peherstorfer, F., On Tchebycheff polynomials on disjoint intervals, in: Haar Memorial

Conference, Vol. II (J. Szabados and K. Tandori, Eds.), North-Holland, Amsterdam, 1987, pp. 737–751.

Petrushev, P.P., & V.A. Popov, Rational Approximation of Real Functions, Cambridge

University Press, Cambridge, 1987. Pinkus, A., On L^{1}-Approximation, Cambridge University Press, Cambridge, 1989. Pinkus, A., & Z. Ziegler, Interlacing properties of the zeros of the error functions in best

L^{p}-approximations, J. Approx. Theory 27 (1979), 1–18.

Pólya, G., Beitrag zur Verallgemeinerung des Verzerrungssatzes auf mehrfach Zusammenhängende Gebiete, S.B. Preuss. Akad. Wiss., Berlin, K.L. Math. Phys. Tech. (1928), 228–232 and 280–282.

Pólya, G., & G. Szegő, Problems and Theorems in Analysis, Vols. I and II, 4th ed., Springer-Verlag, New York, NY, 1976.

Pommerenke, Ch., On some problems by Erdős, Herzog and Piranian, Michigan Math. J. 6 (1959a), 221–225.

Pommerenke, Ch., On some metric properties of polynomials with real zeros, Michigan Math. J. 6 (1959b), 377–380.

Pommerenke, Ch., On the derivative of a polynomial, Michigan Math. J. 6 (1959c), 373–375.

Pommerenke, Ch., On some metric properties of complex polynomials, Michigan Math. J. 8 (1961), 97–115.

Pommerenke, Ch., On some metric properties of polynomials with real zeros, I and II, Michigan Math. J. 6 (1959), 25–27; and 8 (1961), 49–54.

Pommerenke, Ch., Univalent Functions, Vandenhoeck & Ruprecht, Göttingen, 1975.

Protapov, M.K., Some inequalities for polynomials and their derivatives, Vestnik Moskov. Univ. Ser. I Mat. Mekh. 2 (1960), 10–19 (in Russian).

Rahman, Q.I., On extremal properties of the derivatives of polynomials and rational functions, J. London Math. Soc. 35 (1960), 334–336.

Rahman, Q.I., On extremal properties of the derivatives of polynomials and rational functions, Amer. J. Math. 113 (1991), 169–177.

Rahman, Q.I., & G. Labelle, L’influence des zéros sur la borne dans le théoréme de Markoff pour les polynômes sur l’intervalle unité, J. London Math. Soc. 43 (1968), 183–185.

Rahman, Q.I., & G. Schmeisser, Some inequalities for polynomials with a prescribed zero, Trans. Amer. Math. Soc. 216 (1976), 91–103.

Rahman, Q.I., & G. Schmeisser, Les Inégalités de Markoff et de Bernstein, Les Presses de L’Université de Montreal, 1983.

Rahman, Q.I., & G. Schmeisser, L^{p} inequalities for entire functions of exponential type, Trans. Amer. Math. Soc. 320 (1990), 91–103.

Rainville, E.D., Special Functions, Chelsea, New York, NY, 1971.

Rassias, G.M., J.M. Rassias, & Th.M. Rassias, A counter-example to a conjecture of P. Erdős, Proc. Japan Acad. Sci. 53 (1977), 119–121.

Remez, E.J., Sur une propriété des polyômes de Tchebyscheff, Comm. Inst. Sci. Kharkow 13 (1936), 93–95.

Bibliography 463

Reznick, B., An inequality for products of polynomials, Proc. Amer. Math. Soc. 117

(1993). Riesz, M., Eine trigonometrische Interpolations-formel und einige Ungleichungen für

Polynome, Jahresbericht des Deutschen Mathematiker-Vereinigung 23 (1914), 354– 368.

Ritt, J.F., Permutable rational functions, Trans. Amer. Math. Soc. 25 (1923), 399–448. Rivlin, T.J., Chebyshev Polynomials, 2nd ed., Wiley, New York, NY, 1990. Rogers, L.J., An extension of a certain theorem in inequalities, Messenger of Math. 17

(1888), 145–150. Rogosinski, W.W., Extremum problems for polynomials and trigonometric polynomials,

J. London Math. Soc. 29 (1954), 259–275. Royden, H.L., Real Analysis, 3rd ed., MacMillan, New York, NY, 1988. Rubinstein, Z., On a problem of Ilyeff, Pacific J. Math. 26 (1968), 159–161. Rudin, W., Functional Analysis, McGraw-Hill, New York, NY, 1973. Rudin, W., Real and Complex Analysis, 3rd ed., McGraw-Hill, New York, NY, 1987. Saff, E.B., & T. Sheil-Small, Coefficient and integral mean estimates for algebraic and

trigonometric polynomials with restricted zeros, J. London Math. Soc. 9 (1974), 16–22.

Saff, E.B., & V. Totik, Logarithmic Potentials with External Fields, Springer-Verlag (to

appear). Saff, E.B., & R.S. Varga, Uniform approximation by incomplete polynomials, Internat.

J. Math. Math. Sci. 1 (1978), 407–420. Saff, E.B., & R.S. Varga, On lacunary incomplete polynomials, Math. Z. 177 (1981),

297–314. Salem, R., Algebraic Numbers and Fourier Analysis, Heath, Boston, MA, 1963. Schaeffer, A.C., & R.J. Duffin, On some inequalities of S. Bernstein and V. Markoff for

derivatives of polynomials, Bull. Amer. Math. Soc. 44 (1938), 289–297. Scheick, J.T., Inequalities for derivatives of polynomials of special type, J. Approx. The-

ory 6 (1972), 354–358. Schmidt, E., Zur Kompaktheit der Exponentialsummen, J. Approx. Theory 3 (1970),

445–459. Schönhage, A., Asymptotically fast algorithms for the numerical multiplication and di-

vision of polynomials with complex coefficients, in: EUROCAM (1982) Marseille, Springer Lecture Notes in Computer Science, vol. 144, 1982, pp. 3–15.

Schur, I., Über das Maximum des absoluten Betrages eines Polynoms in einen gegebenen Intervall, Math. Z. 4 (1919), 271–287.

464 Bibliography

Schwartz, L., Etude des Sommes d’Exponentielles, Hermann, Paris, 1959. Shapiro, H.S., Topics in Approximation Theory, Springer-Verlag, Berlin, 1971. Smale, S., On the efficiency of the algorithms of analysis, Bull. Amer. Math. Soc. 13

(1985), 87–121. Smith, P.W., An improvement theorem for Descartes systems, Proc. Amer. Math. Soc.

70 (1978), 26–30. Somorjai, G., A Müntz-type problem for rational approximation, Acta. Math. Hungar.

27 (1976), 197–199. Stahl, H., & V. Totik, General Orthogonal Polynomials, Encyclopedia of Mathematics,

vol. 43, Cambridge University Press, New York, NY, 1992. Stewart, I., Galois Theory, Chapman and Hall, London, 1973. Stieltjes, T.J., Oevres Complétes, vol. 2, Groningen, P. Noordhoff, 1914. Stromberg, K.R., Introduction to Classical Real Analysis, Wadsworth, Belmont, CA,

1981. Szabados, J., Bernstein and Markov type estimates for the derivative of a polynomial

with real zeros, in: Functional Analysis and Approximation, Birkhäuser, Basel, 1981, pp. 177–188.

Szabados, J., & A.K. Varma, Inequalities for derivatives of polynomials having real zeros, in: Approximation Theory III (E. W. Cheney, Ed.), Academic Press, New York, NY, 1980, pp. 881–888.

Szabados, J., & P. Vértesi, Interpolation of Functions, World Scientific, Singapore, 1990. Szász, O., Über die Approximation stetiger Funktionen durch lineare Aggregate von

Potenzen, Math. Ann. 77 (1916), 482–496. Szegő, G., Bemerkungen zu einen Satz von J. H. Grace über die Wurzeln algebraischer

Gleighungen, Math. Z. 13 (1922), 28–55. Szegő, G., Über einen Satz von A. Markoff, Math. Z. 23 (1925), 45–61. Szegő, G., Über einen Satz des Herrn Serge Bernstein, Schriften der Königsberger

Gelehrten Gesellschaft 22 (1928), 59–70. Szegő, G., On some problems of approximation, Magyar Tud. Akad. Mat. Kut. Int. Közl.

2 (1964), 3–9. Szegő, G., Orthogonal Polynomials, 4th ed., Amer. Math. Soc., Providence, RI, 1975. Szegő, G., Collected Papers, 3 Vols., Ed. R. Askey, Birkhäuser, Boston MA, 1982. Taslakyan, A.K., Some properties of Legendre quasi-polynomials with respect to a Müntz

system, Mathematics, Erevan University, Erevan (Russian, Armenian summary) 2 (1984), 179-189.

Todd, J., A Legacy from E.I. Zolotarev (1847–1878), Math. Intelligencer 10 (1988), 50–53.

Totik, V., Weighted Approximation with Varying Weight, Lecture Notes in Mathematics, vol. 1569, Springer-Verlag, Berlin, 1994.

Traub, J.F., Iterative Methods for the Solutions of Equations, Chelsea, New York, NY, 1982.

Trent, T.T., A Müntz-Szász theorem for C(D), Proc. Amer. Math. Soc. 83 (1981), 296– 298.

Tsuji, M., Potential Theory in Modern Function Theory, Dover, New York, NY, 1959.

Turán, P., Über die Ableitung von Polynomen, Compositio Math. 7 (1939), 89–95.

Turán, P., On an inequality of C̆ebys̆ev, Ann. Univ. Sci. Budapest. Eötvös Sect. Math. 9 (1968), 15–16.

Turán, P., On a New Method of Analysis and Its Applications, Wiley, New York, NY, 1984.

Vâjâitu, V., & A. Zaharescu, Ilyeff’s conjecture on a corona, Bull. London. Math. Soc. 25 (1993), 49–54.

van Assche, W., & I. Vanherwegen, Quadrature formulas based on rational interpolation, Math. Comp. 61 (1993), 765–783.

van der Poorten, A., A proof that Euler missed... Apéry’s proof of the irrationality of ζ(3), Math. Intelligencer 1 (1979), 195–203.

van der Waerden, B.L., Modern Algebra, Volume II, Frederick Ungar, New York, NY, 1950.

Varga, R., Scientific Computation on Mathematical Problems and Conjectures, SIAM, Philadelphia, PA, 1990.

Varga, R.S., & Wu Wen-da, On the rate of overconvergence of the generalized Eneström-Kakeya functional for polynomials, J. Comput. Math. 3 (1985), 275–288.

Veidinger, L., On the numerical determination of the best approximation in the Chebyshev sense, Numer. Math. 2 (1960), 99–105.

Videnskii, V.S., On estimates of the derivative of a polynomial, Izv. Akad. Nauk SSSR Ser. Mat. 15 (1951), 401–420.

Videnskii, V.S., Markov and Bernstein type inequalities for derivatives of trigonometric polynomials on an interval shorter than the period, Dokl. Akad. Nauk SSSR 130 (1960), 13–16 (in Russian).

Walker, P., Separation of the zeros of polynomials, Amer. Math. Monthly 100 (1993), 272–273.

Walsh, J.L., On the location of the roots of the derivative of a rational function, Ann. of Math. 22 (1920), 128–144.

466 Bibliography

Walsh, J.L., On the location of the roots of the Jacobian of two binary forms, and the

derivative of a rational function, Trans. Amer. Math. Soc. 22 (1921), 101–116. Walsh, J.L., The Location of Critical Points of Analytic and Harmonic Functions, Amer.

Math. Soc., New York, NY, 1950. Walsh, J.L., Interpolation and Approximation by Rational Functions in the Complex

Domain, 5th ed., Amer. Math. Soc., Providence, RI, 1969. Weierstrass, K., Mathematische Werke, Berlin, 1915. Werner, W., On the Simultaneous Determination of Polynomial Roots, Lecture Notes

in Mathematics, vol. 953, Spinger-Verlag, Berlin, 1982, pp. 188–202. Wilf, H.S., Algorithms and Complexity, Prentice Hall, Englewood Cliffs, NJ, 1986. Zhou, S.P., Rational Approximation and Müntz Approximation, Ph. D. Thesis, Dalhousie

University, 1992a. Zhou, S.P., Some remarks on Turán’s inequality, J. Approx. Theory 68 (1992b), 45–48. Zielke, R., Discontinuous Chebyshev Systems, Lecture Notes in Mathematics, vol. 707,

Spinger-Verlag, Berlin, 1979. Zwick, D., Characterization of WT-spaces whose derivatives form a WT-space, J. Ap-

prox. Theory 38 (1983), 188–191. Zygmund, A., Trigonometric Series, Vols. I and II, 2nd ed., Cambridge Univ. Press,

Cambridge, 1977.